OpenStudy (anonymous):

Round to the nearest cent when necessary. Which of the following is the best price? A. 25 for $36.10 B. 30 for $40.00 C. 10 for $23.95 D. 50 for $69.88

OpenStudy (anonymous):

So you need to divide the price by the amount you are getting, to get the price per item. so 36.1/25=1.44

OpenStudy (anonymous):

40/30=1.33 23.95/10=2.395=2.4 69.88/50=1.3976=1.40

OpenStudy (anonymous):

So your answer is B because the cheapest price per item is $1.33